### v3

# Ill-posedness and inviscid limit of the basic equations of fluid dynamics in Besov spaces

Abstract
In this paper, we consider the Cauchy problem to the basic equations of fluid dynamics on the torus. Firstly, we construct a new initial data and provide a simple proof on the ill-posedness of $B^{s}_{p,\infty}$ solution of the Euler equations and the surface quasi-geostrophic equation, which covers the results obtained by Cheskidov-Shvydkoy [10] and Misiołek-Yoneda [29]. Secondly, we prove the failure of the $B^{s}_{p,\infty}$-convergence in the inviscid limit for both the Navier-Stokes equations and the surface quasi-geostrophic equation.
Keywords: Euler and Navier-Stokes equations; QG equation; Ill-posedness; Inviscid limit; Besov spaces
MSC (2020): 35B30; 35Q30; 76U05; 76B03.
1 Introduction
In this paper, we consider the Cauchy problem for the incompressible Navier-Stokes equations in $\mathbb{T}^{d}$ with $d\geq 2$
$\begin{cases}\partial_{t}u+u\cdot\nabla u-\varepsilon\Delta u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{T}^{d},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{T}^{d},\\ u(0,x)=u_{0}(x),&\quad x\in\mathbb{T}^{d}=(\mathbb{R}\setminus 2\pi\mathbb{Z})^{d},\end{cases}$ (NS)
where $\varepsilon>0$ is the kinematic viscosity, the vector field $u(t,x):[0,\infty)\times{\mathbb{T}}^{d}\to{\mathbb{R}}^{d}$ stands for the velocity of the fluid, the quantity $P(t,x):[0,\infty)\times{\mathbb{T}}^{d}\to{\mathbb{R}}$ denotes the scalar pressure, and $\mathrm{div\,}u=0$ means that the fluid is incompressible.
When the viscosity vanishes ($\varepsilon=0$), then the Navier-Stokes equations (NS) reduces to the Euler equations for ideal incompressible fluid
$\begin{cases}\partial_{t}u+u\cdot\nabla u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{T}^{d},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{T}^{d},\\ u(0,x)=u_{0}(x),&\quad x\in\mathbb{T}^{d}=(\mathbb{R}\setminus 2\pi\mathbb{Z})^{d},\end{cases}$ (E)
We say that the Cauchy problem (NS) or (E) is Hadamard (locally) well-posed in a Banach space $X$ if for any data $u_{0}\in X$ there exists (at least for a short time) $T>0$ and a unique solution in the space $\mathcal{C}([0,T),X)$ which depends continuously on the data. In particular, we say that the solution map is continuous if for any $u_{0}\in X$, there exists a neighborhood $B\subset X$ of $u_{0}$ such that for every $u\in B$ the map $u\mapsto U$ from $B$ to $\mathcal{C}([0,T];X)$ is continuous, where $U$ denotes the solution to (NS) or (E) with initial data $u_{0}$. For the well-posedness of (E) in Besov spaces we refer to see [1, Theorem 7.1] and [21, Theorem 1.1].
Next, we mainly recall some of the recent progress which are closely related to our problem. Kato [24] obtained the local well-posedness of classical solution to Euler equations in the Sobolev space $H^{s}(\mathbb{R}^{3})$ for all $s>5/2$. Kato-Ponce [25] extended this result to the Sobolev spaces $W^{s,p}(\mathbb{R}^{3})$ of the fractional order for $s>3/p+1,1<p<\infty$. Chae [6, 7, 8] and Chen-Miao-Zhang [9] gave further extensions to the Triebel-Lizorkin spaces $F_{p,r}^{s}(\mathbb{R}^{3})$ with $s>3/p+1,1<p,r<\infty$ and the Besov spaces $B_{p,r}^{s}(\mathbb{R}^{3})$ with $s>3/p+1$, $1<p<\infty,1\leq r\leq\infty$ or $s=3/p+1,1<p<\infty,r=1$. However, these two kinds of function spaces are only in the $L^{p}(1<p<\infty)$-framework since the Riesz transform is not bounded on $L^{\infty}$. The currently-known best result on the local existence was given by Pak-Park [31] in the Besov space $B_{\infty,1}^{1}(\mathbb{R}^{3})$. Guo-Li-Yin [21] proved the continuous dependence of the Euler equations in the space $B_{p,r}^{s}(\mathbb{R}^{3})$ with $s>3/p+1$, $1\leq p\leq\infty,1\leq r<\infty$ or $s=3/p+1,1\leq p\leq\infty,r=1$. Cheskidov-Shvydkoy [10] proved that the solution of the Euler equations cannot be continuous as a function of the time variable at $t=0$ in the spaces $B^{s}_{p,\infty}(\mathbb{T}^{d})$ where $s>0$ if $2<p\leq\infty$ and s>d(2/p1)s>d(2/p-1) if $1\leq p\leq 2$. Bourgain-Li in [3, 4] employed a combination of Lagrangian and Eulerian techniques to obtain strong local ill-posedness results in borderline Besov spaces $B^{d/p+1}_{p,r}$ for $1\leq p<\infty$ and $1<r\leq\infty$ when $d=2,3$. Misiołek-Yoneda [30] showed that the solution map for the Euler equations is not even continuous in the space of Hölder continuous functions and thus not locally Hadamard well-posed in $C^{1,s}=B^{1+s}_{\infty,\infty}$ with any $s\in(0,1)$.
In this paper, we consider the ill-posedness problem of the Euler equations in Besov spaces. The first result of this paper reads as follows:
Theorem 1.1.
Let $d\geq 2$ and $s>0$ with $1\leq p\leq\infty$. There exists a divergence-free vector field $u_{0}\in B^{s}_{p,\infty}(\mathbb{T}^{d})$ such that the data-to-solution map $u_{0}\mapsto u^{\rm E}(t,u_{0})\in B^{s}_{p,\infty}(\mathbb{T}^{d})$ of the Euler equations (E) satisfies
$\limsup_{t\to 0^{+}}\left\|u^{\rm E}(t,u_{0})-u_{0}\right\|_{B^{s}_{p,\infty}(\mathbb{T}^{d})}\geq\eta_{0},$
where $\eta_{0}$ is some positive constant only dependent on $p$ and $d$.
Remark 1.1.
Theorem 1.1 demonstrates the ill-posedness of the Euler equations in $B^{s}_{p,\infty}$. More precisely, there exists $u_{0}\in B^{s}_{p,\infty}$ such that the corresponding solution to the Euler equations that starts from $u_{0}$ does not converge back to $u_{0}$ in the metric of $B^{s}_{p,\infty}$ as time goes to zero. Theorem 1.1 improves the result obtained by Cheskidov and Shvydkoy in [10] since the index $p>2$ has been enlarged to $1\leq p\leq\infty$. We simplify their approach and present a short and direct proof by constructing explicitly travelling wave solutions, which does not depend on the nonlinear structure of the Euler equations. In fact, we can construct special solutions causing the ill-posedness. To clearly see this, for instance, let $s>0$ and $n$ be a sufficiently large number, we can verify that
$v(x_{1}-t)=2^{-ns}\cos\left(\frac{11}{8}2^{n}(x_{1}-t)\right),\quad v(x_{1})=2^{-ns}\cos\left(\frac{11}{8}2^{n}x_{1}\right)$
are two high frequency wave. In this case, both $\left\|v(x_{1}-t)\right\|_{B_{p,\infty}^{s}(\mathbb{T}^{d})}$ and $\left\|v(x_{1})\right\|_{B_{p,\infty}^{s}(\mathbb{T}^{d})}$ have positive lower and supper bounds. However, notice that
$v(x_{1}-t)=-v(x_{1}),\quad\text{if}\quad\frac{11}{8}2^{n}t=\pi,$
thus the Besov norm of the difference, i.e., $\left\|v(x_{1}-t)-v(x_{1})\right\|_{B_{p,\infty}^{s}(\mathbb{T}^{d})}=2\left\|v(x_{1})\right\|_{B_{p,\infty}^{s}(\mathbb{T}^{d})}$, still has a positive lower bound.
We recall the classical well-posedness theory for the Euler equations, as founded in [1], which is convenient for our purposes: Given the divergence-free initial data $u_{0}\in B_{p,r}^{s}$ with $1\leq p,r\leq\infty$ and $s>1+d/p$, then there exists a short time $T=T\left(u_{0}\right)$, such that the Euler equations (E) has a unique solution $u(t,x)\in C\left([0,T];B_{p,r}^{s}\right)$ if $1\leq r<\infty$ or $u(t,x)\in C_{w}\left([0,T];B_{p,\infty}^{s}\right)$(the subscript $w$ indicates weak continuity in the time variable). Furthermore, we have the estimate
$\|u(t)\|_{B_{p,q}^{s}}\lesssim\left\|u_{0}\right\|_{B_{p,q}^{s}}\quad\text{ for }0\leq t\leq T.$
Our second result of this paper reads as follows:
Theorem 1.2.
Let $d\geq 2$. Assume that $(s,p,r)$ satisfies
$s>\frac{d}{p}+1,(p,r)\in[1,\infty]\times[1,\infty)\quad\mathrm{or}\quad s=\frac{d}{p}+1,(p,r)\in[1,\infty]\times\{1\}.$
For any $\alpha\in(0,1)$, there exists a divergence-free vector field $u_{0}\in B^{s}_{p,r}(\mathbb{T}^{d})$ such that the data-to-solution map $u_{0}\mapsto u^{\rm E}(t,u_{0})\in B^{s}_{p,r}(\mathbb{T}^{d})$ of the Euler equations (E) satisfies
$\limsup_{t\to 0^{+}}\frac{\left\|u^{\rm E}(t,u_{0})-u_{0}\right\|_{B^{s}_{p,r}(\mathbb{T}^{d})}}{t^{\alpha}}=+\infty.$
Remark 1.2.
We would like to mention that Theorem 1.2 is new. In fact, it is known that the solution $u^{\rm E}(t,u_{0})$ for (E) is continuous in time in Besov spaces $B^{s}_{p,r}$ with $r<\infty$, while Theorem 1.2 furthermore indicates that the solution $u^{\rm E}(t,u_{0})$ for (E) cannot be Hölder continuous in time in the same Besov spaces $B^{s}_{p,r}(\mathbb{T}^{d})$.
A classical problem in fluid mechanics is the approximation in the limit $\varepsilon\to 0$ of vanishing viscosity (also called inviscid limit) of solutions of the Euler equations by solutions of the incompressible Navier-Stokes equations. The problem of the convergence of smooth viscous solutions of (NS) to the Eulerian one as $\varepsilon\to 0$ is well understood and has been studied in many literatures, see for example [13, 32, 24], and [15, 16] for the inviscid limit in a bounded domain. Majda [27] showed that under the assumption $u_{0}\in H^{s}$ with $s>\frac{d}{2}+2$, the solutions $u_{\varepsilon}$ to (NS) converge in $L^{2}$ norm as $\varepsilon\to 0$ to the unique solution of Euler equations and the convergence rate is of order $(\varepsilon t)^{\frac{1}{2}}$. Masmoudi [28] improved the result and obtained the convergence in $H^{s}$-norm under the assumption $u_{0}\in H^{s}$ with $s>\frac{d}{2}+1$. In dimension two, Hmidi and Kerrani in [22] proved that (NS) is globally well-posed in Besov space $B^{2}_{2,1}$, with uniform bounds on the viscosity and obtained that the convergence rate of the inviscid limit is of order $\varepsilon t$ for vanishing viscosity. Subsequently, in [23], they further generalized to other Besov spaces $B^{2/p+1}_{p,1}$ with convergence in $L^{p}$. Chemin [5] resolved inviscid limit of Yudovich type solutions with only the assumption that the vorticity is bounded. In the case of $\mathbb{T}^{2}$ or $\mathbb{R}^{2}$, by taking greater advantage of vorticity formulation, more beautiful results were obtained quantitatively (see for example [2, 11, 12] and the references therein). Guo-Li-Yin [21] solved the inviscid limit in the same topology. However, it left an open problem for the end-point case $r=\infty$ in the inviscid limit of the Navier-Stokes equations in Besov spaces. Our third result is the following.
Theorem 1.3.
(Non-convergence) Let $d\geq 2$ and $\varepsilon\in[0,1]$. Assume that $s>1+d/p$ with $1\leq p\leq\infty$. Then a family of solution maps $(u_{0},\varepsilon)\mapsto u^{\rm NS}_{\varepsilon}(t,u_{0})$ to the Navier-Stokes equations (NS) do not converge to the solution map $u_{0}\mapsto u^{\rm E}(t,u_{0})$ of the Euler equations (E) in $B^{s}_{p,\infty}$. More precisely, there exists an initial data $u_{0}\in U_{R}$ such that
$\limsup_{\varepsilon_{n}\to 0^{+}}\left\|u^{\rm NS}_{\varepsilon_{n}}(t_{n},u_{0})-u^{\rm E}(t_{n},u_{0})\right\|_{B^{s}_{p,\infty}(\mathbb{T}^{d})}\geq\eta_{0},$
with some positive constant $\eta_{0}$ only dependent on $p$ and $d$.
Next, we consider the Cauchy problem for the two-dimensional surface quasi-geostrophic equation which is a fundamental example of active scalar transport
$\begin{cases}\partial_{t}\theta+u\cdot\nabla\theta+\kappa\Lambda^{\alpha}\theta=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{T}^{2},\\ u=\mathcal{R}^{\perp}\theta:=\nabla^{\perp}\Lambda^{-1}\theta,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{T}^{2},\\ \theta(0,x)=\theta_{0}(x),&\quad x\in\mathbb{T}^{2}=(\mathbb{R}\setminus 2\pi\mathbb{Z})^{2},\end{cases}$ (QG)
where the unknown $\theta=\theta(x,t)$ is scalar field, $\Lambda^{\alpha}f$ is defined via the Fourier modes of $f$
$\widehat{\Lambda^{\alpha}f}(k)=|k|^{\alpha}\widehat{f}(k).$
$\mathcal{R}=\left(\mathcal{R}_{1},\mathcal{R}_{2}\right)$ is the vector of Riesz transforms, then
$u=\left(\mathcal{R}_{2}\theta,-\mathcal{R}_{1}\theta\right),\quad\widehat{\mathcal{R}_{j}\theta}(\xi)=-i\frac{\xi_{j}}{|\xi|}\hat{\theta}(\xi),\quad\xi_{j}\in\mathbb{Z}\quad j=1,2.$
When the viscosity vanishes ($\kappa=0$), (QG) reduces to the inviscid quasi-geostrophic equation. This equation was introduced in [14] as a two-dimensional model of the 3D Euler equations. There have been significant development in the ill-posedness theory (see e.g.,[17, 18, 20, 33]) for the QG equation. To put our study in the proper perspective, we first recall a few results which are related to our problem.
Ill-posedness in $W^{1,\infty}$: Elgindi and Masmoudi [19] proved the ill-posedness of active scalar system in $L^{\infty}$-type spaces in the sense that there exist smooth steady states $\bar{\theta}$ and a sequence of perturbations $\tilde{\theta}_{0}^{(\epsilon)}\left(\epsilon\rightarrow 0^{+}\right)$ such that the associated solution $\theta^{(\epsilon)}$ with data $\bar{\theta}+\tilde{\theta}_{0}^{(\epsilon)}$ satisfies
$\left\|\theta^{(\epsilon)}(0,\cdot)-\bar{\theta}\right\|_{W^{1,\infty}}<\epsilon,\quad\sup_{0<t<\epsilon}\left\|\theta^{(\epsilon)}(t,\cdot)-\bar{\theta}\right\|_{W^{1,\infty}}>c(\bar{\theta}).$
Ill-posedness in $B_{p,\infty}^{s}$: Kiselev, Nazarov and Volberg [26] observed that if $\theta_{0}\in L^{p}(\mathbb{T}^{2})$ with $1<p<\infty$ then the solution of (QG) with $\alpha=1$ satisfies $\lim_{t\rightarrow 0^{+}}\left\|\theta(t)-\theta_{0}\right\|_{L^{p}}=0$. Misiołek and Yoneda [29] showed that this property fails in certain Besov spaces for $0<\alpha<1$ by using elementary properties of certain lacunary Fourier series.
Theorem 1.4 ([29]).
Let $0<\alpha<1$. There exists $\theta_{0}\in B_{p,\infty}^{s}(\mathbb{T}^{2})$ such that the corresponding (weak) solution $\theta$ of the Cauchy problem (QG) satisfies
$\lim_{t\rightarrow 0^{+}}\left\|\theta(t)-\theta_{0}\right\|_{B_{p,\infty}^{s}(\mathbb{T}^{2})}>0$
for any $s>0$ and any $2\leq p\leq\infty$.
Our aim is to generalize the above result to the broader cases.
Theorem 1.5 ($\kappa>0$).
Let $s>0,\alpha\in(0,2]$ and $1\leq p\leq\infty$. There exists an initial data $\theta_{0}\in B^{s}_{p,\infty}(\mathbb{T}^{2})$ such that the data-to-solution map $\theta_{0}\mapsto\theta(t)\in B^{s}_{p,\infty}(\mathbb{T}^{2})$ of the surface quasi-geostrophic equation (QG) satisfies for some positive constant $\eta_{0}$ which is only dependent on $p$
$\limsup_{t\to 0^{+}}\left\|\theta(t)-\theta_{0}\right\|_{B^{s}_{p,\infty}(\mathbb{T}^{2})}\geq\eta_{0}.$
Remark 1.3.
Theorem 1.5 demonstrates the ill-posedness of the QG equations in $B^{s}_{p,\infty}$. More precisely, there exists $u_{0}\in B^{s}_{p,\infty}$ such that the corresponding solution to the QG equations that starts from $u_{0}$ does not converge back to $u_{0}$ in the metric of $B^{s}_{p,\infty}$ as time goes to zero. We should mention that Theorem 1.5 holds for the generalised QG equation (see e.g.,[17, 18])
$\begin{cases}\partial_{t}\theta+u\cdot\nabla\theta+\kappa\Lambda^{\alpha}\theta=0,\\ u=\nabla^{\bot}\Lambda^{\beta-2}\theta,\\ \theta(0,x)=\theta_{0}(x).\end{cases}$
Theorem 1.6.
(Non-convergence) Let $\alpha\in(0,2]$ and $s>1+\frac{2}{p}$ with $1\leq p\leq\infty$. There exists an initial data $\theta_{0}\in B^{s}_{p,\infty}(\mathbb{T}^{2})$ such that the solution $\theta^{\kappa}(t)$ of the dissipative (QG) does not converge to the solution $\theta^{0}(t)$ of the inviscid (QG) for small $t\in(0,T_{1}]$ in $B^{s}_{p,\infty}$ as $\kappa\downarrow 0$. More precisely, there exists an initial data $\theta_{0}\in B^{s}_{p,\infty}(\mathbb{T}^{2})$ such that for a short time $t=\kappa$
$\limsup_{\kappa\to 0^{+}}\left\|\theta^{\kappa}(t,\theta_{0})-\theta^{0}(t,\theta_{0})\right\|_{B^{s}_{p,\infty}(\mathbb{T}^{2})}\geq\eta_{0},$
with some positive constant $\eta_{0}$ depending on $p$ and $\delta$ but independent on $\kappa$.
Theorem 1.7 ($\kappa=0$).
Let $s>0$ and $1\leq p\leq\infty$. There exists an initial data $\theta_{0}\in B^{s}_{p,\infty}(\mathbb{T}^{2})$ such that the data-to-solution map $\theta_{0}\mapsto\theta(t)\in B^{s}_{p,\infty}(\mathbb{T}^{2})$ of the inviscid (QG) satisfies for some positive constant $\eta_{0}$ which is only dependent on $p$
$\limsup_{t\to 0^{+}}\left\|\theta(t)-\theta_{0}\right\|_{B^{s}_{p,\infty}\cap L_{\mathcal{R}}^{\infty}(\mathbb{T}^{2})}\geq\eta_{0},$ (1.1)
where we denote
$\left\|f\right\|_{B^{s}_{p,\infty}\cap L_{\mathcal{R}}^{\infty}(\mathbb{T}^{2})}:=\left\|f\right\|_{B^{s}_{p,\infty}(\mathbb{T}^{2})}+\left\|f\right\|_{L^{\infty}(\mathbb{T}^{2})}+\|\mathcal{R}^{\perp}f\|_{L^{\infty}(\mathbb{T}^{2})}.$
Remark 1.4.
Theorem 1.7 implies the ill-posedness of the inviscid (QG) either in $B^{s}_{p,\infty}$ or in $L^{\infty}$.
Corollary 1.1 ($\kappa=0$).
Let $s>\frac{2}{p}$ and $1\leq p<\infty$. There exists an initial data $\theta_{0}\in B^{s}_{p,\infty}(\mathbb{T}^{2})$ such that the data-to-solution map $\theta_{0}\mapsto\theta(t)\in B^{s}_{p,\infty}(\mathbb{T}^{2})$ of the inviscid (QG) satisfies for some positive constant $\eta_{0}$ which is only dependent on $p$
$\limsup_{t\to 0^{+}}\left\|\theta(t)-\theta_{0}\right\|_{B^{s}_{p,\infty}(\mathbb{T}^{2})}\geq\eta_{0}.$
Remark 1.5.
The comparison between the proof of Theorem 1.5 and Theorem 1.7 tells us that the mechanics between the dissipative and inviscid QG equation leading to the discontinuous of data-to-solution at zero in the weaker Besov spaces is completely different. Precisely speaking, the primarily affect which leads to the ill-posedness of the dissipative QG equation is the diffusion term while for the inviscid QG equation it is the convection term.
2 Preliminaries
We define the periodic Fourier transform $\mathcal{F}_{\mathbb{T}^{d}}:\mathcal{D}(\mathbb{T}^{d})\rightarrow\mathcal{S}(\mathbb{Z}^{d})$ as
$(\mathcal{F}u)(k)=\widehat{u}(k)=\frac{1}{(2\pi)^{d}}\int_{\mathbb{T}^{d}}e^{-\mathrm{i}x\cdot k}u(x)\mathrm{d}x.$
We decompose $u\in\mathcal{D}(\mathbb{T}^{d})$ on the circle $\mathbb{T}^{d}$ into Fourier series, i.e.
$u(x)=\sum_{k\in\mathbb{Z}^{d}}\widehat{u}(k)e^{\mathrm{i}x\cdot k}.$
We are interested in solutions which take values in the Besov space $B_{p,r}^{s}(\mathbb{T}^{d})$. Recall that one way to define this space requires a dyadic partition of unity. Given a smooth bump function $\chi$ supported on the ball of radius $4/3$, and equal to 1 on the ball of radius $3/4$, we set $\varphi(\xi)=\chi(2^{-1}\xi)-\chi(\xi)$ and $\varphi_{j}(\xi)=\varphi(2^{-j}\xi)$ and then we deduce that $\varphi$ satisfies that ${\rm{supp}}\;\varphi\subset\left\{\xi\in\mathbb{T}^{d}:3/4\leq|\xi|\leq 8/3\right\}$ and $\varphi(\xi)\equiv 1$ for $4/3\leq|\xi|\leq 3/2$. Using this partition, we define the periodic dyadic blocks as follows
$\Delta_{j}u=0,\quad\text{ if }\quad j\leq-2,$
$\Delta_{-1}u=\sum_{\xi\in\mathbb{Z}^{d}}\chi(\xi)\widehat{u}(\xi)e^{\mathrm{i}x\cdot\xi},$
$\Delta_{j}u=\sum_{\xi\in\mathbb{Z}^{d}}\varphi_{j}(\xi)\widehat{u}(\xi)e^{\mathrm{i}x\cdot\xi},\quad\text{ if }\quad j\geq 0.$
The operators $\Delta_{j}$ defined on the periodic domain share many properties with those on the whole space(see [1]). In particular, we obtain the Littlewood-Paley decomposition of $u$
$u=\sum_{j\geq-1}\Delta_{j}u\quad\text{in}\quad\mathcal{S}^{\prime}(\mathbb{T}^{d}).$
Definition 2.1.
Let $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. The nonhomogeneous Besov space $B^{s}_{p,r}(\mathbb{T}^{d})$ is defined by
$B^{s}_{p,r}(\mathbb{T}^{d}):=\left\{f\in\mathcal{S}^{\prime}(\mathbb{T}^{d}):\;\|f\|_{B^{s}_{p,r}(\mathbb{T}^{d})}<\infty\right\},$
where
$\|f\|_{B^{s}_{p,r}(\mathbb{T}^{d})}=$ $\left(\sum_{j\geq-1}\left(2^{sj}\|\Delta_{j}f\|_{L^{p}(\mathbb{T}^{d})}\right)^{r}\right)^{1/r},$ if $1\leq r<\infty$,
$\|f\|_{B^{s}_{p,r}(\mathbb{T}^{d})}=$ $\sup_{j\geq-1}\left(2^{sj}\|\Delta_{j}f\|_{L^{p}(\mathbb{T}^{d})}\right),$ if $r=\infty$.
Let us complete this section by presenting two lemmas which will be used often in the sequel.
Lemma 2.1.
Let $3\leq m\in\mathbb{Z}$ and $1\leq j\in\mathbb{Z}$, we have
$\Delta_{j}\left[\cos\left(\frac{11}{8}2^{m}x_{1}\right)\right]=\begin{cases}0,&\text{if}\quad j\neq m,\\ \cos\left(\frac{11}{8}2^{m}x_{1}\right),&\text{if}\quad j=m.\end{cases}$
Proof.
We set $\lambda=\frac{11}{8}2^{m}$ and $\vec{\lambda}=(\lambda,0,\cdots,0)$ for simplicity. Notice that $\varphi_{j}(k)$ is symmetric, i.e., $\varphi_{j}(k)=\varphi_{j}(|k|)$, we deduce
$\Delta_{j}\left(\cos(\lambda x_{1})\right)$ $=\frac{1}{2}\sum_{k\in\mathbb{Z}^{d}}\varphi_{j}(k)\left(\mathbf{1}_{\vec{\lambda}}(k)+\mathbf{1}_{-\vec{\lambda}}(k)\right)e^{\mathrm{i}x\cdot k}$
$=\frac{1}{2}\sum_{\ell\in\mathbb{Z}}\varphi_{j}(\ell)\left(\mathbf{1}_{\lambda}(\ell)+\mathbf{1}_{-\lambda}(\ell)\right)e^{\mathrm{i}x_{1}\ell}$
$=\varphi_{j}(\lambda)\cos(\lambda x_{1})=\begin{cases}0,&\text{if}\quad j\neq m,\\ \cos(\lambda x_{1}),&\text{if}\quad j=m.\end{cases}$
where $\mathbf{1}_{K}(x)$ is the indicator function, taking a value of 1 if $x=K$ and 0 otherwise.
Thus we obtain the desired result of Lemma 2.1. ∎
Remark 2.1.
For large fixed $m$, $\varphi_{j}\left(\frac{11}{8}2^{m}x_{1}\right)$ vanish except for $j=m$ due to the support condition of $\varphi$ and $\varphi(k)\equiv 1$ for $4/3\leq|k|\leq 3/2$.
The following simple fact is needed in the sequel.
Lemma 2.2.
Let $1\ll n\in\mathbb{Z}$, we have
$\left\|\cos\left(\frac{11}{8}2^{n}x\right)\right\|_{L^{p}([0,2\pi])}=c_{0}:=\begin{cases}\left(2\int^{\pi}_{0}|\cos x|^{p}\mathrm{d}x\right)^{1/{p}},&\text{if}\quad p\in[1,\infty),\\ 1,&\text{if}\quad p=\infty.\end{cases}$
3 Proof of Theorem 1.1
Choice of initial data. Let us fix an $s>0$ and define the initial data
$u_{0}(x)=(1,f(x_{1}),0,\ldots,0),$ (3.1)
where $f(x)$ is a bounded real-valued periodic function of one variable with the following form
$f(x)=\sum\limits^{\infty}_{j=3}2^{-js}\cos\left(\frac{11}{8}2^{j}x\right).$ (3.2)
It is not difficult to check that $\mathord{{\rm div}}u_{0}=0$ and $u_{0}\in B^{s}_{p,\infty}(\mathbb{T}^{d})$ for any $1\leq p\leq\infty$.
Construction of solution. An fundamental observe is that the vector field
$u(t,x)=\left(1,f\left(x_{1}-t\right),0,\ldots,0\right)$ (3.3)
is an obvious periodic solution of the incompressible Euler equations:
$\partial_{t}u+u\cdot\nabla u+\nabla P=0,\quad\mathrm{div\,}u=0,$
with $P=0$, i.e. this is a pressureless flow.
From (3.1) and (3.3), we have
$u(t,x)-u_{0}(x)=\left(0,f\left(x_{1}-t\right)-f(x_{1}),0,\ldots,0\right),$
From the explicit formula for $f(x)$ in (3.2) and by direct computation, we see that
$f(x_{1}-t)-f(x_{1})$ $=\sum\limits^{\infty}_{j=3}2^{-js}\left[\cos\left(\frac{11}{8}2^{j}x_{1}\right)\mathbf{a}_{j}(t)+\sin\left(\frac{11}{8}2^{j}x_{1}\right)\mathbf{b}_{j}(t)\right],$
where
$\mathbf{a}_{j}(t)=\cos\left(\frac{11}{8}2^{j}t\right)-1\quad\text{and}\quad\mathbf{b}_{j}(t)=\sin\left(\frac{11}{8}2^{j}t\right).$ (3.4)
Then, by Lemma 2.1, we have for some $n$ large enough
$2^{ns}\Delta_{n}[f(x_{1}-t)-f(x_{1})]$ $=\cos\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{a}_{n}(t)+\sin\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{b}_{n}(t).$
Letting $\frac{11}{8}2^{n}t_{n}=\pi$, then one has $\mathbf{a}_{n}(t_{n})=-2$ and $\mathbf{b}_{n}(t_{n})=0$. Thus we have for some $n$ large enough
$\quad 2^{ns}\left\|\Delta_{n}[f(x_{1}-t_{n})-f(x_{1})]\right\|_{L^{p}(\mathbb{T}^{d})}=2(2\pi)^{d-1}\left\|\cos\left(\frac{11}{8}2^{n}x_{1}\right)\right\|_{L^{p}(\mathbb{T})}.$
From which and Lemma 2.2, we deduce that for some $n$ large enough
$\|u(t_{n},x)-u_{0}(x)\|_{B^{s}_{p,\infty}(\mathbb{T}^{d})}$ $=\|f(x_{1}-t_{n})-f(x_{1})\|_{B^{s}_{p,\infty}(\mathbb{T}^{d})}$
$\geq 2^{ns}\left\|\Delta_{n}[f(x_{1}-t_{n})-f(x_{1})]\right\|_{L^{p}(\mathbb{T}^{d})}$
$=2c_{0}(2\pi)^{d-1}.$
Notice that $t_{n}\to 0^{+}$ as $n\to\infty$, we complete the proof of Theorem 1.1.
4 Proof of Theorem 1.2
Let us fix an $s>1+\frac{d}{p}$ and define the initial data
$u_{0}(x)=(1,g(x_{1}),0,\cdots,0),$
where $g(x)$ is a bounded real-valued period function of one variable with the following form
$g(x)=\sum\limits^{\infty}_{j=3}j^{-2}2^{-js}\cos\left(\frac{11}{8}2^{j}x\right).$ (4.1)
It is not difficult to check that $\mathord{{\rm div}}u_{0}=0$ and $u_{0}\in B^{s}_{p,r}(\mathbb{T}^{d})$ for any $1\leq p,r\leq\infty$.
Let $t\geq 0$, we consider
$u(t,x)=\left(1,g\left(x_{1}-t\right),0,\ldots,0\right),$
Obviously, it is a period function which also satisfies the incompressible Euler equations (E) with initial condition $u_{0}(x)=(1,g(x_{1}),0,\cdots,0)$.
Continue in a similar fashion we see that
$u(t,x)-u_{0}(x)=\left(0,g\left(x_{1}-t\right)-g(x_{1}),0,\ldots,0\right),$
and
$g(x_{1}-t)-g(x_{1})$ $=\sum\limits^{\infty}_{j=3}j^{-2}2^{-js}\left(\cos\left(\frac{11}{8}2^{j}x_{1}\right)\mathbf{a_{j}}(t)+\sin\left(\frac{11}{8}2^{j}x_{1}\right)\mathbf{b}_{j}(t)\right),$
where $\mathbf{a_{j}}(t)$ and $\mathbf{b_{j}}(t)$ are given by (3.4).
Then, by Lemma 2.1, we have for some $n$ large enough
$n^{2}2^{ns}\Delta_{n}[g(x_{1}-t)-g(x_{1})]$ $=\cos\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{a}_{n}(t)+\sin\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{b}_{n}(t),$
Letting $\frac{11}{8}2^{n}t_{n}=\pi$, then one has $\mathbf{a}_{n}(t_{n})=-2$ and $\mathbf{b}_{n}(t_{n})=0$. Thus we have for some $n$ large enough
$\quad 2^{ns}\left\|\Delta_{n}[g(x_{1}-t_{n})-g(x_{1})]\right\|_{L^{p}(\mathbb{T}^{d})}=2(2\pi)^{d-1}n^{-2}\left\|\cos\left(\frac{11}{8}2^{n}x_{1}\right)\right\|_{L^{p}(\mathbb{T})}.$
From which and Lemma 2.2, we deduce that for some $n$ large enough
$t_{n}^{-\alpha}\|u(t_{n},x)-u_{0}(x)\|_{B^{s}_{p,r}(\mathbb{T}^{d})}$ $=t_{n}^{-\alpha}\|g(x_{1}-t_{n})-g(x_{1})\|_{B^{s}_{p,r}(\mathbb{T}^{d})}$
$\geq t_{n}^{-\alpha}2^{ns}\left\|\Delta_{n}[g(x_{1}-t_{n})-g(x_{1})]\right\|_{L^{p}(\mathbb{T}^{d})}$
$=2c_{0}(2\pi)^{d-1}t_{n}^{-\alpha}n^{-2}.$
Notice that $t_{n}\to 0^{+}$ and $t_{n}^{-\alpha}n^{-2}\to+\infty$ as $n\to\infty$, we complete the proof of Theorem 1.2.
5 Proof of Theorem 1.3
Choice of initial data. Let us fix an $s$ satisfying $s>1+d/p$ with $1\leq p\leq\infty$ and define the initial data
$u_{0}(x)=(1,f(x_{1}),0,\ldots,0),$
where $f(x)$ is given by (3.2).
Construction of solution to the Euler equations.
Let $t\geq 0$, we consider
$u^{\rm E}(t,x)=\left(1,f(x_{1}-t),0,\ldots,0\right).$
By the classical well-posedness theory, we know that $u^{\rm E}(t,x)$ is a unique periodic solution of the Euler equations (E) with initial data $u_{0}(x)=\left(1,f(x_{1}),0,\ldots,0\right)$.
Construction of solution to the Navier-Stokes equations.
Assume that $f_{n}(t,x_{1})$ solves the Cauchy Problem
$\begin{cases}\partial_{t}f_{n}+\partial_{x_{1}}f_{n}-\varepsilon_{n}\partial^{2}_{x_{1}}f_{n}=0,\quad t>0,\\ f_{n}(t=0,x_{1})=f(x_{1}).\end{cases}$ (5.1)
Let $t\geq 0$, we consider
$u^{\rm NS}_{\varepsilon_{n}}(t,x)=(1,f_{n}(t,x_{1}),0,\ldots,0).$
Due to (5.1), we know that $u^{\rm NS}_{n}(t,x)$ is a unique periodic solution of the Navier-Stokes equations with initial data $u_{0}(x)=\left(1,f(x_{1}),0,\ldots,0\right)$.
It is easy to deduce that the Cauchy Problem (5.1) has a unique explicit solution
$f_{n}(t,x_{1})=\sum\limits^{\infty}_{j=3}2^{-js}\left[\cos\left(\frac{11}{8}2^{j}(x_{1}-t)\right)e^{-\varepsilon_{n}\frac{121}{64}2^{2j}t}\right].$
We shall compare the solution of Euler equations with that of Navier-Stokes equations. Obviously,
$u^{\rm NS}_{\varepsilon_{n}}(t_{n},u_{0})-u^{\rm E}(t_{n},u_{0})=(0,f_{n}(t,x_{1})-f(x_{1}-t_{n}),0,\ldots,0),$
and
$f_{n}(t,x_{1})-f(x_{1}-t)$ $=\sum\limits^{\infty}_{j=3}2^{-js}\cos\left(\frac{11}{8}2^{j}(x_{1}-t)\right)\left[e^{-\varepsilon_{n}\frac{121}{64}2^{2j}t}-1\right].$
By Lemma 2.1, one has for some $n$ large enough
$\Delta_{n}\left[f_{n}(t,x_{1})-f(x_{1}-t_{n})\right]=2^{-ns}\left[\cos\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{e}_{n}(t)+\sin\left(\frac{11}{8}2^{n}x_{1}\right)\mathbf{f}_{n}(t)\right]\mathbf{g}_{n}(t),$
where
$\mathbf{e}_{n}(t)=\cos\left(\frac{11}{8}2^{n}t\right),\quad\mathbf{f}_{n}(t)=\sin\left(\frac{11}{8}2^{n}t\right)\quad\text{and}$
$\mathbf{g}_{n}(t)=e^{-\varepsilon_{n}\frac{121}{64}2^{2n}t}-1.$
Letting $\frac{11}{8}2^{n}t_{n}=\pi$ and $\varepsilon_{n}=\frac{8}{11\pi}2^{-n}$, which gives that
$\left(\mathbf{e}_{n}(t_{n}),\;\mathbf{f}_{n}(t_{n}),\;\mathbf{g}_{n}(t_{n})\right)=(-1,\;0,\;e^{-1})$
and thus
$\Delta_{n}\left[f_{n}(t_{n},x_{1})-f(x_{1}-t_{n})\right]$ $=(1-e^{-1})2^{-ns}\cos\left(\frac{11}{8}2^{n}x_{1}\right).$
Combining the above, we have
$\|u^{\rm NS}_{\varepsilon_{n}}(t_{n},x)-u^{\rm E}(t_{n},u_{0})\|_{B^{s}_{p,\infty}(\mathbb{T}^{d})}$ $\geq 2^{ns}\left\|\Delta_{n}\left(f_{n}(t_{n},x_{1})-f(x_{1}-t_{n})\right)\right\|_{L^{p}(\mathbb{T}^{d})}$
$=\left(1-e^{-1}\right)\cdot\left\|\cos\left(\frac{11}{8}2^{n}x_{1}\right)\right\|_{L^{p}(\mathbb{T}^{d})}$
$=c_{0}\left(1-e^{-1}\right)(2\pi)^{d-1}.$
Taking the $\limsup$, we deduce that
$\limsup_{\varepsilon_{n}\to 0^{+}}\left\|u^{\rm NS}_{\varepsilon_{n}}(t_{n},u_{0})-u^{\rm E}(t_{n},u_{0})\right\|_{B^{s}_{p,\infty}}\geq c_{0}(2\pi)^{d-1}.$
Notice that $t_{n}\to 0^{+}$ and $\varepsilon_{n}\to 0^{+}$ as $n\to\infty$, we complete the proof of Theorem 1.3.
6 Proof of Theorems 1.5-1.6
Choice of initial data. Let us fix an $s>0$ and define the initial data which is a bounded real-valued periodic function of one variable with the following form
$\theta_{0}(\mathbf{x})=f(x_{1})=\sum\limits^{\infty}_{j=3}2^{-js}\cos\left(\frac{11}{8}2^{j}x_{1}\right),\quad\mathbf{x}=(x_{1},x_{2})\in\mathbb{T}^{2}.$
It is not difficult to check that $\theta_{0}\in B^{s}_{p,\infty}(\mathbb{T}^{2})$ for any $1\leq p\leq\infty$.
Construction of solution. We are interested in solutions of the form
$\theta(t,\mathbf{x})=f\left(t,x_{1}\right),$
then it must have
$u\cdot\nabla\theta=u_{1}\partial_{1}\theta+u_{2}\partial_{2}\theta=0.$
In fact,
$\partial_{2}\theta=0\quad\text{and}\quad u_{1}=-\Lambda^{-1}\partial_{2}\theta=0.$
Thus $\theta(t,\mathbf{x})=e^{-\kappa\Lambda^{\alpha}t}\theta_{0}(x_{1})$ is a periodic solution of (QG). We shall compare the solution of (QG) with the initial data. Obviously,
$\theta(t,\mathbf{x})-\theta_{0}(\mathbf{x})$ $=\left(e^{-\kappa\Lambda^{\alpha}t}-1\right)\theta_{0}(\mathbf{x}).$
By Lemma 2.1, one has for some $n$ large enough
$\Delta_{j}\left(\left(e^{-\kappa\Lambda^{\alpha}t}-1\right)\cos\left(\frac{11}{8}2^{n}x_{1}\right)\right)$ $=\begin{cases}0,&\text{if}\quad j\neq n,\\ \left(e^{-\kappa\left(\frac{11}{8}2^{n}\right)^{\alpha}t}-1\right)\cos\left(\frac{11}{8}2^{n}x_{1}\right),&\text{if}\quad j=n,\end{cases}$
which gives that
$\Delta_{n}\left[\theta(t,\mathbf{x})-\theta_{0}(\mathbf{x})\right]=2^{-ns}\cos\left(\frac{11}{8}2^{n}x_{1}\right)\left[e^{-\kappa\left(\frac{11}{8}2^{n}\right)^{\alpha}t}-1\right].$ (6.1)
Proof of Theorem 1.5. Letting $\kappa\left(\frac{11}{8}2^{n}\right)^{\alpha}t_{n}=1$ in (6.1), one has
$\Delta_{n}\left[\theta(t_{n},\mathbf{x})-\theta_{0}(\mathbf{x})\right]$ $=(1-e^{-1})2^{-ns}\cos\left(\frac{11}{8}2^{n}x_{1}\right).$
Combining the above, we have
$\|\theta(t_{n},\mathbf{x})-\theta_{0}(\mathbf{x})\|_{B^{s}_{p,\infty}(\mathbb{T}^{2})}$ $\geq 2^{ns}\left\|\Delta_{n}\left[\theta(t_{n},\mathbf{x})-\theta_{0}(\mathbf{x})\right]\right\|_{L^{p}(\mathbb{T}^{2})}$
$=\left(1-e^{-1}\right)\cdot\left\|\cos\left(\frac{11}{8}2^{n}x_{1}\right)\right\|_{L^{p}(\mathbb{T}^{2})}$
$=2\pi c_{0}\left(1-e^{-1}\right).$
Taking the $\limsup$, we deduce that
$\limsup_{t_{n}\to 0^{+}}\left\|\theta(t_{n},\mathbf{x})-\theta_{0}\right\|_{B^{s}_{p,\infty}}\geq\pi c_{0}.$
Thus we complete the proof of Theorem 1.5.
Proof of Theorem 1.6. We should notice that, the (QG) with $\kappa=0$ has a unique steady-state solution $\theta^{0}(t,\mathbf{x})=\theta_{0}(\mathbf{x})$. Let $\kappa_{n}=t_{n}=\left(\frac{11}{8}2^{n}\right)^{-\frac{\alpha}{2}}$ in (6.1), which gives that
$\Delta_{n}\left[\theta^{\kappa_{n}}(t_{n},\mathbf{x})-\theta^{0}(t_{n},\mathbf{x})\right]$ $=(1-e^{-1})2^{-ns}\cos\left(\frac{11}{8}2^{n}x_{1}\right),$
and thus we deduce that
$\limsup_{\kappa_{n}\to 0^{+}}\left\|\theta^{\kappa_{n}}(t_{n},\mathbf{x})-\theta^{0}(t_{n},\mathbf{x})\right\|_{B^{s}_{p,\infty}}\geq\pi c_{0}.$
Notice that $t_{n}\to 0^{+}$ and $\kappa_{n}\to 0^{+}$ as $n\to\infty$, we complete the proof of Theorem 1.6.
7 Proof of Theorem 1.7
Lemma 7.1.
Let $s>0$. We define an initial data $\theta_{0}$ as follows
$\theta_{0}(\mathbf{x}):=\sum^{\infty}_{j=3}2^{-js}\sin\left(\frac{11}{8}2^{j}x_{1}\right)+\sin x_{2},\quad\mathbf{x}=(x_{1},x_{2})\in\mathbb{T}^{2}.$
Then there exists some sufficiently large $n\in\mathbb{Z}^{+}$ and some positive constants $C,c$ such that
$\left\|\theta_{0}\right\|_{B^{s}_{p,\infty}\cap L^{p}(\mathbb{T}^{2})}\leq C,$
$2^{ns}\left\|\Delta_{n}(u_{0}\cdot\nabla\theta_{0})\right\|_{L^{p}(\mathbb{T}^{2})}\geq c2^{n}.$
Proof.
The first is obvious. Notice that
$\Lambda^{-1}\cos(\lambda x_{i})=\frac{1}{2}\sum_{k\in\mathbb{Z}^{2}}\frac{1}{|k|}\left(\mathbf{1}_{\vec{\lambda}}(k)+\mathbf{1}_{-\vec{\lambda}}(k)\right)e^{\mathrm{i}x\cdot k}=\frac{1}{\lambda}\cos(\lambda x_{i}),\quad i=1,2,$
then
$u_{0}^{(1)}(x):=-\Lambda^{-1}\partial_{x_{2}}\theta_{0}(\mathbf{x})=-\cos x_{2},$
$u_{0}^{(2)}(x):=\Lambda^{-1}\partial_{x_{1}}\theta_{0}(\mathbf{x})=\sum^{\infty}_{j=3}2^{-js}\cos\left(\frac{11}{8}2^{j}x_{1}\right),$
and thus
$u_{0}\cdot\nabla\theta_{0}=\cos x_{2}\cdot\sum^{\infty}_{j=3}2^{-js}\left(1-\frac{11}{8}2^{j}\right)\cos\left(\frac{11}{8}2^{j}x_{1}\right).$
From which, we have
$\Delta_{n}(u_{0}\cdot\nabla\theta_{0})=\cos x_{2}\cdot 2^{-ns}\left(1-\frac{11}{8}2^{n}\right)\cos\left(\frac{11}{8}2^{n}x_{1}\right),$
which in turn gives
$2^{ns}\left\|\Delta_{n}(u_{0}\cdot\nabla\theta_{0})\right\|_{L^{p}(\mathbb{T}^{2})}$ $=\left(\frac{11}{8}2^{n}-1\right)\left\|\cos x_{2}\right\|_{L^{p}([0,2\pi])}\left\|\cos\left(\frac{11}{8}2^{n}x_{1}\right)\right\|_{L^{p}([0,2\pi])}$
$=\left(\frac{11}{8}2^{n}-1\right)\left(2\int^{\pi}_{0}|\cos x|^{p}\mathrm{d}x\right)^{2/{p}}\geq c2^{n},$
thus we complete the proof of Lemma 7.1. ∎
The following Lemma involves the nonlinear perturbation for the solution map of (QG) and is standard.
Lemma 7.2.
Assume that $s>0$ and $1\leq p\leq\infty$. Then we have for $t\in(0,T]$
$\left\|u\theta-u_{0}\theta_{0}\right\|_{\dot{B}^{s}_{p,\infty}}\leq C\left(\|\theta-\theta_{0}\|_{B^{s}_{p,\infty}}\|\theta_{0}\|_{L^{\infty}}+\|\theta-\theta_{0}\|_{L_{\mathcal{R}}^{\infty}}\|\theta_{0}\|_{B^{s}_{p,\infty}}\right).$ (7.1)
From (QG) and the Newton-Leibniz formula, it follows that
$\theta(t)-\theta_{0}$ $=-tu_{0}\cdot\nabla\theta_{0}-\int_{0}^{t}\mathord{{\rm div}}\left(u\theta-u_{0}\theta_{0}\right)\mathrm{d}\tau.$
Using the triangle inequality and Lemma 7.2, we have
$\left\|\theta(t)-\theta_{0}\right\|_{B^{s}_{p,\infty}}$ $\geq 2^{{ns}}\left\|\Delta_{n}\left(\theta(t)-\theta_{0}\right)\right\|_{L^{p}}$
$\geq t2^{{n}s}\left\|\Delta_{n}\big(u_{0}\cdot\nabla\theta_{0}\big)\right\|_{L^{p}}-t2^{n}2^{{n}s}\left\|\Delta_{n}(u\theta-u_{0}\theta_{0})\right\|_{L^{p}}$
$\geq t2^{n}\left(c-C\|\theta_{0}\|_{B^{s}_{p,\infty}\cap L^{\infty}}\|\theta(t)-\theta_{0}\|_{B^{s}_{p,\infty}\cap L_{\mathcal{R}}^{\infty}}\right).$
Suppose that (1.1) were not true for some $s>0$ and $1\leq p\leq\infty$. We pick $0<t_{0}\leq t$ such that for all $\tau\in[0,t_{0}]$
$\left\|\theta(t)-\theta_{0}\right\|_{B^{s}_{p,\infty}\cap L_{\mathcal{R}}^{\infty}}<\min\left(1,\frac{c}{2C\|\theta_{0}\|_{B^{s}_{p,\infty}\cap L^{\infty}}}\right).$
Thus, picking $t_{n}=M2^{-n}$ with large $M>0$, we have
$1>\left\|\theta(t_{n})-\theta_{0}\right\|_{B^{s}_{p,\infty}}\geq\frac{c}{2}M,$
which leads to a contraction for large $M$. This completes the proof of Theorem 1.7.
Declarations
Data Availability
No data was used for the research described in the article.
Conflict of interest
The authors declare that they have no conflict of interest.
Funding
J. Li is supported by the National Natural Science Foundation of China (12161004), Innovative High end Talent Project in Ganpo Talent Program (gpyc20240069), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (20232BCJ23009), Jiangxi Provincial Natural Science Foundation (20252BAB22004). X.Wu is supported by the Natural Science Foundation of Henan Province (252300421774).